# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.3.1 When $\mathrm{IE}(\sigma_{*}^{I})=2$ with $I\geq 2$
Starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $C_{2}d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}\kappa+\eta\tilde{\eta}\nu,$ (B.16)
where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
Proof.  We first compute one gradient step from $\boldsymbol{w}=\boldsymbol{\omega}$ with a fresh sample $(\boldsymbol{x},y)$.
~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{x}.$ (B.17)
Then, with a projection matrix $\boldsymbol{P}_{\boldsymbol{\omega}}=I-\boldsymbol{\omega}\boldsymbol{\omega}^{\top}$, the updated parameter becomes
$\boldsymbol{w}\leftarrow\boldsymbol{w}+\boldsymbol{P}_{\boldsymbol{\omega}}\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x},$ (B.18)
and the next gradient step with the same sample is computed as
~$\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ $=\eta y\sigma^{\prime}(\boldsymbol{w}^{\top}\boldsymbol{x})\boldsymbol{x}$ (B.19)
$=y\sigma^{\prime}\big{(}(\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})^{\top}\boldsymbol{x}\big{)}\boldsymbol{x}$ (B.20)
$=y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{x},$ (B.21)
here we used the notation $\|\boldsymbol{v}\|_{\boldsymbol{A}}^{2}=\boldsymbol{v}^{\top}\boldsymbol{A}\boldsymbol{v}$ for a vector $\boldsymbol{v}\in\mathbb{R}^{d}$ and a positive symmetric matrix $\boldsymbol{A}\in\mathbb{R}^{d\times d}$. From (B.18) and (B.21), the parameter after the two steps is obtained as
$\boldsymbol{w}$ $\leftarrow\boldsymbol{w}+\tilde{\nabla}_{\boldsymbol{w}}y\sigma(\boldsymbol{w}^{\top}\boldsymbol{x})$ (B.22)
$=\boldsymbol{\omega}+\eta y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+\eta y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}$ (B.23)
$=\boldsymbol{\omega}+\eta\boldsymbol{g},$ (B.24)
where we defined
$\boldsymbol{g}:=y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}.$ (B.25)
Finally, normalization yields
$\boldsymbol{w}$ $\leftarrow\frac{\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})}{\|\boldsymbol{w}-\xi(\boldsymbol{w}-\boldsymbol{\omega})\|}=\frac{\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.26)
Therefore, the update of the alignment is
$\boldsymbol{\theta}^{\top}\boldsymbol{w}$ $=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{(1+\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2})^{\frac{1}{2}}}$ (B.27)
$\geq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.28)
On the other hand, we have
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}.$ (B.29)
We evaluate the expectation of (B.28). For the $j$-th Hermite polynomial $\mathsf{He}_{j}$ and $\boldsymbol{u}\in\mathbb{S}^{d-1}$, we have that
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{i}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{1}^{\top}\boldsymbol{x}]$ (B.30)
$=j({\sf e}_{1}^{\top}x)^{j-1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]+({\sf e}_{1}^{\top}\boldsymbol{x})^{j+1}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})],$ (B.31)
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}(x_{1})f(\boldsymbol{u}^{\top}\boldsymbol{x}){\sf e}_{2}^{\top}\boldsymbol{x}]=({\sf e}_{1}^{\top}\boldsymbol{x})^{j}({\sf e}_{2}^{\top}\boldsymbol{x})\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})].$ (B.32)
Therefore,
$\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[\mathsf{He}_{j}({\sf e}_{1}^{\top}\boldsymbol{x})f(\boldsymbol{u}^{\top}\boldsymbol{x})\boldsymbol{x}]$ (B.33)
$=\begin{pmatrix}j({\sf e}_{1}^{\top}\boldsymbol{x})^{j-1}\\ 0\\ \vdots\\ 0\end{pmatrix}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j-1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]{\sf e}_{1}+({\sf e}_{1}^{\top}\boldsymbol{x})^{j}\mathbb{E}_{\boldsymbol{x}\sim\mathcal{N}(0,\boldsymbol{I}_{d})}[f^{(j+1)}(\boldsymbol{u}^{\top}\boldsymbol{x})]\boldsymbol{u}.$ (B.34)
Hence the first term of $\boldsymbol{g}$ (B.25) can be exapanded as
$\mathbb{E}[y\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ $=\boldsymbol{P}_{\boldsymbol{\omega}}\mathbb{E}\bigg{[}\bigg{(}\sum_{j=p}^{q}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}j\sum_{j=0}^{C_{q}}\beta_{j}\mathsf{He}_{j}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\bigg{)}\boldsymbol{x}\bigg{]}$ (B.35)
$=\boldsymbol{P}_{\boldsymbol{\omega}}\sum_{j=p}^{q}\bigg{[}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j}\boldsymbol{\omega}\bigg{]}$ (B.36)
$=\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}.$ (B.37)
The coefficient is evaluated as
$\bigg{|}\sum_{j=p}^{q}j!\alpha_{j}\beta_{j}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\big{)}^{j-1}\bigg{|}\lesssim\kappa^{p-1}\leq\kappa^{2}.$ (B.38)
For the second term of $\boldsymbol{g}$ (B.25), we first bound the difference in replacing $\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}$ with $d$,
$\big{|}\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]-\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]\big{|}$ (B.39)
$=2\big{|}\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x})]\big{|},$ (B.40)
where $h$ is a polynomial with degree at most $(C_{q}+q)^{C_{q}-1}+q+1$ and coefficients are all $O(1)$. (B.40) is further upper bounded as
(B.40)$\leq 2\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}\mathbb{E}[(h(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2},\boldsymbol{\omega}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{x},\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}))^{2}]^{\frac{1}{2}},$ (B.41)
by Cauchy-Schwarz inequality. $\mathbb{E}[(\eta\|\boldsymbol{x}\|_{\boldsymbol{P}_{\boldsymbol{\omega}}}^{2}-\eta d)^{2}]^{\frac{1}{2}}=\eta(2d-1)^{\frac{1}{2}}$, and the expectation of $\mathbb{E}[h^{2}]$ is $O(1)$ when $\eta\leq d^{-1}$. Therefore, (B.40) is bounded by $C_{1}\eta d^{\frac{1}{2}}$.
Now, we consider $\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$. The following decomposition can be made.
$\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.42)
$=\sum_{j=1}^{C_{\sigma}}j\beta_{j}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.43)
$=\sum_{k=0}^{j-1}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose k}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.44)
We evaluate each term of (B.44) except for $k=I-1$. Each term of (B.44) is a constant multiple of $\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{\theta}$. and we can evaluate the constant by
$\bigg{|}\boldsymbol{\theta}^{\top}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}^{k}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.45)
$=(\eta d)^{k}\bigg{|}\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}^{k+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}.$ (B.46)
When $k\leq I-2$, $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})^{k+1}$ has information exponent larger than $2$. Therefore, we have

### v2

@@ B.4 Weak Recovery: Stochastic Update
Take $\eta^{2t},\eta^{2t+1}=\eta=c_{\eta}d^{-1}$, $\xi^{2(t+1)}=\xi=1-c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$. Suppose that the link function satisfies $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}$ (we choose the smallest such $I$) and activation functions satisfy all of the assumptions in Section B.1 for weak recovery. Let
$T_{1,1}=C_{3}c_{\xi}^{-1}\begin{cases}d&(\text{if $p_{*}=\mathrm{GE}(\sigma_{*})=1$})\\ d(\log d)&(\text{else if $p_{*}=\mathrm{GE}(\sigma_{*})=2$})\\ d^{p_{*}-1}&(\text{else $p_{*}=\mathrm{GE}(\sigma_{*})\geq 3$}),\end{cases}$ (B.76)
and take $c_{\xi}\lesssim\delta\mathrm{poly}(c_{\eta})$, $c_{2}\gtrsim\mathrm{poly}(c_{\eta})$, and $C_{3}\simeq c_{1}^{-1}$. If $\kappa^{0}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$, there exists some $\tau_{*}\leq T_{1,1}$ such that
$\kappa^{2\tau_{*}}\geq 2c_{2},$ (B.77)
with probability at least $1-\delta$, and $\kappa^{2\tau}\geq 2c_{2}$ for all $\tau_{*}\leq\tau\leq T_{1,1}$, with high probability.
We may take $\delta=o_{d}(1)$ with arbitrarily slow decay. The proof is adapted from [BAGJ21], but our bound on $T_{1,1}$ is slightly sharper (by a $\log d$ factor for $p_{*}=2$ and by a $(\log d)^{2}$ factor for $p_{*}\geq 3$). For $p_{*}=2$, this is because of a trick that we carefully “restart” the dynamics, whose failure probability exponentially decays.
Proof.  We divide the proof into the following cases.
(i) When $p_{*}=1$. Note that $\{\sum_{s=0}^{\tau}\nu^{2s}\}_{\tau}$ is Martingale with $\mathbb{E}[(\nu^{2s})^{2}]\lesssim 1$. By Doob’s maximal inequality and Markov’s inequality, with probability $1-\delta$, we have
$\max_{0\leq\tau\leq T}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}^{2}\leq\delta^{-1}\mathbb{E}[(\sum_{s=0}^{T}\nu^{2s})^{2}]\leq\delta^{-1}\sum_{s=0}^{T}\mathbb{E}[(\nu^{2s})^{2}]\leq C_{1}\delta^{-1}(T+1)$ (B.78)
for any fixed $T\geq 0$, with a sufficiently large constant $C_{1}=\Theta(1)$. In the following we consider the case when (B.78) holds for $T=c_{1}^{-1}c_{\xi}^{-1}d-1$.
If $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}\leq c_{\eta}^{I}$ for all $t=0,1,\cdots,\tau$, we have
$\kappa^{2(\tau+1)}$ $\geq\kappa^{2\tau}+c_{1}c_{\xi}d^{-1}+c_{\eta}c_{\xi}d^{-1}\nu^{2\tau}$ (B.79)
$\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}(\tau+1)d^{-1}\gamma-c_{\eta}c_{\xi}d^{-1}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}.$ (B.80)
Now, applying (B.78) to get